OpenStudy (anonymous):

A function is given. Determine the average rate of change of the function between the given values of the variable. f(z) = 2 − 3z^2; z = −2, z = 0

OpenStudy (anonymous):

average rate of change is just \[ \frac{f(b)-f(a)}{b-a} \]

OpenStudy (anonymous):

In this case \(b=0\) and \(a=-2\)

OpenStudy (shamil98):

so f(0) = 2 - 3(0)^2 = f(b) f(-2) = 2 - 3(-2)^2 = f(a) Evaluate and use the formula wio has given above to determine the average rate of change.